# BadBlocks: Low-Cost and Stealthy Backdoor Attacks Tailored for Text-to-Image Diffusion Models

### v1

@@ 4.1 Attack Configuration
In the experimental section, we adopted standard settings for the backdoor attack. For the text-to-image task, we used three different characters as the backdoor trigger, including invisible Unicode, specific phrase, and lowercase greek letter. As attack targets, we selected three mainstream diffusion models: Stable Diffusion v1.5, Stable Diffusion v2.1 and Realistic Vision v4.0, which were trained on the MS-COCO-Caption dataset[45]. To assess the quality of the generated images, we use the FID score as the primary metric[46]. Furthermore, we employ LPIPS and SSIM to evaluate the attack success rate (ASR)[47, 48]. All training is conducted for 5 epochs, with a learning rate of 1e41e-4 and a batch size of 8. All training and evaluation processes were conducted on an NVIDIA A40 GPU, but this does not mean that professional computing graphics cards are necessary for BadBlocks. Our experimental results demonstrate that consumer-grade GPUs, such as the NVIDIA RTX 3060 (12GB) and RTX 4070Ti SUPER (16GB), can also perform efficient and successful backdoor attacks within a short time frame. Table.1 presents a comparison between BadBlocks and existing approaches in terms of parameter modification scope and GPU resource consumption.
4.2 Attack Results
Table.2 presents the attack results of BadBlocks on three different baseline models. The experimental results prove that establishing a backdoor mapping requires only a few key weights. In Stable Diffusion v1.5, the FID loss was merely 4.3%, whereas the UNet-based method exhibited a loss of 73.4% under the same poisoning rate. In the other two baselines, BadBlocks also demonstrated similar effects. The experimental results demonstrate the feasibility of implementing a fine-grained backdoor training scheme by decoupling the model architecture, indicating that backdoor attacks often require infection of only the key components. In Appendix.C, we present the attack results under different schedulers.
@@ 4.2 Attack Results
Table 2: We evaluated the performance of BadBlocks across three different baseline models and found that it not only executed backdoor attacks effectively, but also exhibited strong stealthiness.
4.3 Advantages and Limitations
Experimental results demonstrate that backdoor attacks based on diffusion models only require minimal parameter training. Effective attacks can be achieved by updating a small subset of model parameters. The primary advantage of our method lies in its ability to significantly reduce parameter modifications, thereby lowering computational costs and reducing training time. Furthermore, BadBlocks is highly compatible with other attack strategies, as its effectiveness does not rely on loss function manipulation. In Appendix.E, we combine BadBlocks with mainstream attack methods, significantly reducing resource consumption without compromising attack performance and enhancing concealment. Despite the demonstrated efficiency of BadBlocks, certain limitations remain. Experimental results indicate that the poisoning rate (PR) significantly affects the quality of the target images generated. When the PR falls below 10%, the target model is still capable of producing backdoor images. However, because of the limited reconstruction capacity of a single sampled block, it is not capable of generating high-resolution targets. This constraint imposes certain limitations on the types of datasets that attackers can effectively exploit. Figure.5 shows the attack results of BadBlocks under different poisoning rates in three baseline models.
Figure 6: A single component is insufficient for backdoor mapping, effective attacks require at least the ResNet and Transformer blocks.
5 Ablation Study
@@
5.1 Impact of Crucial Components
To explore component-level roles in backdoor attacks and reduce the attack surface, we conducted ablation studies on the ResNet layer, Transformer block, and normalization layer within the sampling block. Results show that single-component fine-tuning fails to inject a backdoor, while jointly attacking the normalization and Transformer blocks enables backdoor injection at the cost of degraded image quality. As shown in Figure.6, although the ResNet layer is not directly involved in backdoor expression, it is essential for preserving normal generation quality.
5.2 Different Number of Infected Blocks
In general, the upsampling stage of a diffusion model consists of multiple sequentially connected upsampling blocks. We conducted ablation studies by varying the number of infected sample blocks used for fine-tuning. The experimental results reveal that increasing the number of fine-tuned blocks does not necessarily enhance backdoor performance. In contrast, an excessive number of infected blocks may hinder effective backdoor mapping, suppress backdoor activation, and induce excessive assimilation effects. Figure.7 shows model performance with backdoor injection across different numbers of injected blocks.
@@ Appendix A Technical Appendix
Appendix B Evaluating BadBlocks via L-norms
@@
Appendix C The variation of parameter distribution
@@
Appendix D Performance under different schedulers
@@
Appendix E Visualized Attack Results
@@
Appendix F Optimize Existing work with BadBlocks
@@ F.1 RickRolling
Table 3: It is evident that integrating BadBlocks can significantly enhance attack efficiency and generation quality, while further improving resistance to attention-based detection methods.
In Table.3, after applying BadBlocks, RickRolling’s image generation quality increased by up to 38.3%, reduced BDR by up to 90.6%, and cut GPU time by up to 62.5%. This is sufficient to demonstrate the advantage of BadBlocks in terms of attack performance.
F.2 BadDiffusion
BadDiffusion is a well-established attack method within the image input space that activates the backdoor mapping by embedding a patch trigger into the image. Building upon BadDiffusion, we have developed an algorithm based on BadBlocks specifically for image-to-image tasks. Details of the algorithm can be found in Algorithm.2. We will discuss the attack performance in Table.4.
@@ F.2 BadDiffusion
Table 4: Since BadDiffusion relies on patch-based triggers, BadBlocks cannot reduce the visual prominence of these patches, and thus existing defense frameworks targeting BadDiffusion may remain effective. However, this does not negate the reduced FID loss and lower resource consumption achieved by integrating BadBlocks.
F.3 VillanDiffusion
@@
Therefore, in this part, we present detailed results of BadBlocks under various sampling methods, including both ODE-based and SDE-based samplers such as DDIM, DDPM, DPM-Solver, and PNDM. As shown in Table.5, a comparison of attack performance reveals that nearly all schedulers exhibit performance improvements when integrated with BadBlocks.
@@
Table 5: Even under different schedulers, BadBlocks consistently activates backdoors and achieves exceptionally high attack success rates . Obviously, the random noise introduced by SDE-based sampling fails to disrupt the backdoor mapping of BadBlocks, further highlighting its robustness as a persistent threat.

### v4

@@ IV-A Attack Configuration
In the experimental section, we adopted standard settings for the backdoor attack. For the text-to-image task, we used three different characters as the backdoor trigger, including invisible Unicode, specific phrase, and lowercase greek letter. As attack targets, we selected three mainstream diffusion models: Stable Diffusion v1.5, Stable Diffusion v2.1 and Realistic Vision v4.0, which were trained on the MS-COCO-Caption dataset[coco]. To assess the quality of the generated images, we use the FID score as the primary metric[FID]. Furthermore, we employ LPIPS and SSIM to evaluate the attack success rate (ASR)[LPIPS, ssim]. All training is conducted for 5 epochs, with a learning rate of 1e41e-4 and a batch size of 8. All training and evaluation processes were conducted on an NVIDIA A40 GPU, but this does not mean that professional computing graphics cards are necessary for BadBlocks. Our experimental results demonstrate that consumer-grade GPUs, such as the NVIDIA RTX 3060 (12GB) and RTX 4070Ti SUPER (16GB), can also perform efficient and successful backdoor attacks within a short time frame. Table.I presents a comparison between BadBlocks and existing approaches in terms of parameter modification scope and GPU resource consumption.
IV-B Attack Results
Table II reports the attack performance of BadBlocks on three baseline models. The results show that establishing an effective backdoor mapping requires updating only a small subset of key parameters, indicating strong redundancy in diffusion models that can be exploited by adversaries. On Stable Diffusion v1.5, BadBlocks incurs only 4.3% FID degradation, while the UNet-based method suffers a loss of 73.4% under the same poisoning rate. Similar trends are observed on the other two baselines, demonstrating the generality of our method. These results validate the feasibility of fine-grained backdoor training via architectural decoupling, suggesting that infecting only critical components is sufficient to achieve high attack success. Furthermore, by substantially reducing parameter modifications, BadBlocks lowers computational cost, shortens training time, and alleviates GPU memory requirements. Owing to its independence from specific loss manipulations or additional regularization, BadBlocks is highly compatible with existing attack strategies. As shown in Appendix.E, integrating BadBlocks with mainstream attacks significantly reduces resource consumption without sacrificing attack success or stealthiness. Additional results under different schedulers are provided in Appendix.C.
TABLE II: We evaluated the performance of BadBlocks across three different baseline models and found that it not only executed backdoor attacks effectively, but also exhibited strong stealthiness.
@@ IV-B Attack Results
V Ablation Study
@@
V-A Impact of Crucial Components
To explore component-level roles in backdoor attacks and reduce the attack surface, we conducted ablation studies on the ResNet layer, Transformer block, and normalization layer within the sampling block. Results show that single-component fine-tuning fails to inject a backdoor, while jointly attacking the normalization and Transformer blocks enables backdoor injection at the cost of degraded image quality. As shown in Figure 6, although the ResNet layer is not directly involved in backdoor expression, it is essential for preserving normal generation quality.
Figure 6: The activation of the backdoor process cannot be achieved without the participation of key network components. A single component is insufficient for backdoor mapping, we found effective attacks require at least the ResNet and Transformer blocks in upsampling blocks.
V-B Different Number of Infected Blocks
In general, the upsampling stage of a diffusion model consists of multiple sequentially connected upsampling blocks. We conducted ablation studies by varying the number of infected sample blocks used for fine-tuning. The experimental results reveal that increasing the number of fine-tuned blocks does not necessarily enhance backdoor performance. In contrast, an excessive number of infected blocks may hinder effective backdoor mapping, suppress backdoor activation, and induce excessive assimilation effects. Figure 7 shows model performance with backdoor injection across different numbers of injected blocks.
@@ Appendix A Technical Appendix
A-A A.Evaluating BadBlocks via L-norms
@@
A-B B.The variation of parameter distribution
@@
A-C C.Performance under different schedulers
@@
A-D D. Visualized Attack Results
@@
A-E E.Optimize Existing work with BadBlocks
@@ A-E1 RickRolling
TABLE III: It is evident that integrating BadBlocks can significantly enhance attack efficiency and generation quality, while further improving resistance to attention-based detection methods.
In Table.III, after applying BadBlocks, RickRolling’s image generation quality increased by up to 38.3%, reduced BDR by up to 90.6%, and cut GPU time by up to 62.5%. This is sufficient to demonstrate the advantage of BadBlocks in terms of attack performance.
A-E2 BadDiffusion
BadDiffusion is a well-established attack method within the image input space that activates the backdoor mapping by embedding a patch trigger into the image. Building upon BadDiffusion, we have developed an algorithm based on BadBlocks specifically for image-to-image tasks. Details of the algorithm can be found in Algorithm.2. We will discuss the attack performance in Table.IV.
@@ A-E2 BadDiffusion
TABLE IV: Since BadDiffusion relies on patch-based triggers, BadBlocks cannot reduce the visual prominence of these patches, and thus existing defense frameworks targeting BadDiffusion may remain effective. However, this does not negate the reduced FID loss and lower resource consumption achieved by integrating BadBlocks.
A-E3 VillanDiffusion
@@
Therefore, in this part, we present detailed results of BadBlocks under various sampling methods, including both ODE-based and SDE-based samplers such as DDIM, DDPM, DPM-Solver, and PNDM. As shown in Table.V, a comparison of attack performance reveals that nearly all schedulers exhibit performance improvements when integrated with BadBlocks.
@@
TABLE V: Even under different schedulers, BadBlocks consistently activates backdoors and achieves exceptionally high attack success rates . Obviously, the random noise introduced by SDE-based sampling fails to disrupt the backdoor mapping of BadBlocks, further highlighting its robustness as a persistent threat.